OpenStudy (anonymous):

use analytic methods to find the extreme values of h(x)= ln(x+1) on the interval 0≤x≤3

OpenStudy (anonymous):

First check the endpoints: \[f(0)=\ln(0+1)=0\\ f(3)=\ln(3+1)=\ln4\] Now use the derivative test: \[f'(x)=\frac{1}{x+1}\] The derivative is never zero, and over the interval \(0\le x\le3\), the value of the derivative is decreasing. What can you say about the extreme values?